### v1

# An Inverse Problem for the Prescribed Mean Curvature

Abstract.
We extend the study of inverse problems for minimal surfaces by considering the inverse source problem for the prescribed mean curvature equation
$\nabla\cdot\left[\frac{\nabla u}{(1+|\nabla u|^{2})^{1/2}}\right]=H(x).$
We prove that in two dimensions, the source function $H$ is uniquely determined by the associated Dirichlet-to-Neumann map. A notable feature of this problem is that although the equation is posed on an Euclidean domain, its linearization yields an anisotropic conductivity equation where the coefficient matrix corresponds to a Riemannian metric $g$ depending on the background solution. This work represents the first treatments of inverse source problems for quasilinear equations.
The proof relies on the higher order linearization method. The main methodological contribution is the development of an approach to decouple the resulting system of nonlinear algebraic and geometric equations, which enables the complete determination of the source term from boundary measurements. The decoupling uses a Liouville type uniqueness result for conformal mappings.
Keywords. Inverse problems, inverse source problems, prescribed mean curvature equation, quasilinear elliptic equations.
1. Introduction
We consider an inverse problem for the prescribed mean curvature equation for $2$-dimensional surfaces living in $\mathbb{R}^{3}$. The equation in $\Omega\subset\mathbb{R}^{n}$ is
(1.1) $\nabla\cdot\left[\frac{\nabla u}{(1+\lvert\nabla u\rvert^{2})^{1/2}}\right]=H(x),$
where $H\in C^{\alpha}(\Omega)$. Let us assume for now that the Dirichlet problem for (1.1) is well-posed on an open subset $\mathcal{N}\subset C^{2,\alpha}(\partial\Omega)$. In this case, the Dirichlet-to-Neumann map (DN map) is defined by the usual assignment
(1.2) $\Lambda_{H}:\mathcal{N}\to C^{1,\alpha}(\partial\Omega),\qquad f\mapsto\left.\partial_{\nu}u_{f}\right|_{\partial\Omega},$
where $u_{f}$ is the unique (small) solution to (1.1) with boundary value $f$. Here $\nu$ denotes the unit outer normal on $\partial\Omega$.
The well-posedness holds if $H\in C^{1}(\Omega)$ and is small enough when compared to the isoperimetric constant and mean curvature of the boundary of $\Omega$, see [GT01, Theorem 16.10]. By the same theorem for some
$f_{0}\in C^{2,\alpha}(\partial\Omega)\text{ there is a unique solution }u_{0}\in C^{2,\alpha}(\overline{\Omega}).$
Now we can make a small perturbation by $f\in C^{2,\alpha}(\partial\Omega)$ to the boundary value $f_{0}$ to obtain a unique solution $u_{f}$ of (1.1) with $u_{f}|_{\partial\Omega}=f_{0}+f$. Then we can use the implicit function theorem to say that this solution depends smoothly on $f$. Notice that by [GT01, Theorem 16.10] we do not need to restrict ourselves to unique solutions in some neighborhood of $u_{0}$, sometimes called small solutions in previous works. We refer to [LL24, Theorem 2.1] for a related well posedness proof.
We ask the following question:
Inverse problem: Can we determine $H$ from the DN map?
This type of an inverse problem is called an inverse source problem. We will give a positive answer to this:
Theorem 1.1.
Let $H,\tilde{H}\in C^{\infty}(\overline{\Omega})$ be such that $H|_{\partial\Omega}=\tilde{H}|_{\partial\Omega}$ to high order and $f_{0}\in C^{\infty}(\partial\Omega)$. Assume that $u_{0}$ and $\tilde{u}_{0}$ solve
(1.3) $\left\{\begin{array}[]{ll}\nabla\cdot\left[\frac{\nabla u}{(1+\lvert\nabla u\rvert^{2})^{1/2}}\right]=H(x),&\text{in}\,\,\Omega\\ u=f_{0},&\text{on}\,\,\partial\Omega,\end{array}\right$.
for $H$ and $\tilde{H}$ on the right hand side respectively, and with $u_{0}|_{\partial\Omega}=\tilde{u}_{0}|_{\partial\Omega}=f_{0}$. Suppose that the DN maps of (1.1) satisfy
$\Lambda_{H}f=\Lambda_{\tilde{H}}f,\quad\text{for all }f\in C^{\infty}(\partial\Omega)$
with $\lVert f_{0}-f\rVert_{C^{\infty}(\partial\Omega)}\leq\delta$ for sufficiently small $\delta>0$. Then $H=\tilde{H}$ in $\Omega$.
1.1. Motivation
The prescribed mean curvature (PMC) equation is a prototypical quasilinear elliptic equation of the form
(1.4) $\nabla\cdot\left(A(\nabla u)\nabla u\right)=H(x),$
where $A:\mathbb{R}^{n}\to\mathbb{R}$. The main motivation of the present work is to serve as a first step in solving inverse problems for equations of the above. To the best of our knowledge, it is the first to consider an inverse source problem for a quasilinear elliptic equation.
Another key motivation for our study concerns the method we employ to solve the inverse problem. Our approach is based on the higher order linearization technique. The first step of this method involves solving the Calderón problem for the first linearized equation
(1.5) $\begin{cases}\nabla\cdot\left(g^{-1}\nabla v\right)=0,&\text{in }\Omega,\\ v=f,&\text{on }\partial\Omega,\end{cases}$
where $g$ is a positive definite symmetric matrix field, thus a Riemannian metric. This linearized equation is an anisotropic conductivity equation. Therefore, our study provides a new application for the theory of the anisotropic Calderón problem, even though the original nonlinear equation is posed on a Euclidean domain.
Regarding the PMC equation itself, it generalizes the minimal surface equation ($H=0$) by requiring a surface to have a specific, pre-assigned mean curvature $H$ at every point, rather than zero curvature. Physically, this corresponds to modeling surfaces in a state of balanced tension under external forces, such as soap bubbles with internal pressure, rather than area-minimizing soap films. Consequently, the PMC equation serves as a fundamental model in diverse areas, including capillary surfaces, biological membranes, and general relativity, where it arises in the initial data formulation of Einstein’s equations [Sei97, CGP10].
Beyond these classical applications, we mention that inverse problems for minimal surfaces have recently gained surprising significance within the AdS/CFT duality theory in theoretical physics. There, techniques from inverse problems have proven effective in such cases [JLST25] where the mathematical tools traditionally used in physics have been ineffective. Due to this surprising application, we mention the works [HMS11, Don14] that study surfaces with prescribed curvature in holography.
1.2. Earlier works
Our work is a new addition to a growing branch of inverse problems for nonlinear partial differential equations. In [Isa93] the author linearized the DN map to solve an inverse problem for a nonlinear PDE. By linearizing one can then use methods for linear equations to solve the problem. In recent years the interest for these has grown and it started with the works [KLU18, FO20a, LLLS21a]. In these works the method of higher order linearization was introduced. In this method one linearizes the DN map multiple times and solves an inverse problem at each step, gaining new information about the original problem.
This method has been applied to many problems. Quasilinear elliptic isotropic conductivity equations have been considered in [CFK+21, KKU20, LW23]. In these works the goal is to recover the nonlinear conductivity. Also inverse problems for the minimal surface equation, quasilinear in nature, on Riemannian manifolds and Euclidean domains have been considered in [ABN20, CLT24, CLLT23, CLLO24, Nur23b, Nur24], and in the context of AdS/CFT duality in physics in [JLST25]. Also in [MT24] the author studies a Calderón type problem for harmonic maps. A map between two Riemannian manifolds is harmonic if it satisfies a certain quasilinear equation.
Inverse problems for elliptic second order semilinear equations have been considered in [IS94, IN95, Sun10, IY13, LLLS21a, LLLS21b, LLST22, KU20, FO20b, FLL23, ST23, Nur23a, JNS23, JNS25] and inverse problems for nonlinear magnetic Schrödinger equations in [KU23, LZ23, Ma20, KUY24]. We also mention the works for nonlinear biharmonic equations [BKSU25, NS25], $p$-harmonic equation [BKS15, BIK18, CF25] and the works for nonlinear fractional equations [Lin22, LL19a, LL23, LL19b]. See also [Las25] for a recent survey on this subject.
The above mentioned works consider PDEs with zero right hand side, that is, they do not have a source term. Recently in [LL24, KLL24] the authors considered inverse problems for nonlinear partial differential equations where the aim is to recover a nonlinear term and a source term. They also demonstrate by example nonlinearities that one can break a gauge appearing in a general case of recovery.
1.3. Method of proof
The methods we use can be divided into two main parts. First, we combine information from the Dirichlet-to-Neumann (DN) map with the higher-order linearization technique and complex geometric optics (CGO) solutions for the Schrödinger equation to derive equations for quantities related to the background solution $u_{0}$ (corresponding to the boundary value $f_{0}$). These equations are highly coupled, and the second step involves solving for $\nabla u_{0}$ from this system. A key ingredient at this stage is a Liouville-type uniqueness result for conformal mappings.
To show that $H=\tilde{H}$ in $\Omega$, we proceed as follows. If $u_{0}$ and $\tilde{u}_{0}$ are the solutions of the prescribed mean curvature equation corresponding to the sources $H$ and $\tilde{H}$ respectively, and if we establish that $\nabla u_{0}=\nabla\tilde{u}_{0}$, then it follows immediately from the structure of the equation that
$H=\nabla\cdot\left[\frac{\nabla u_{0}}{(1+|\nabla u_{0}|^{2})^{1/2}}\right]=\nabla\cdot\left[\frac{\nabla\tilde{u}_{0}}{(1+|\nabla\tilde{u}_{0}|^{2})^{1/2}}\right]=\tilde{H}$.
We now provide a detailed outline of the argument. The first linearization of (1.1) yields an anisotropic conductivity equation, equivalent to a Schrödinger equation involving the Laplace–Beltrami operator associated with the Riemannian metric whose inverse is
$g^{-1}=\frac{1}{(1+|\nabla u_{0}|^{2})^{1/2}}\left(I_{2\times 2}-\frac{\nabla u_{0}\otimes\nabla u_{0}}{1+|\nabla u_{0}|^{2}}\right).$
This metric explicitly depends on the background solution $u_{0}$. Using boundary information about $H$, we show that the DN maps for the corresponding Schrödinger equations coincide. Applying [CLLO24, Theorem 1.1], we conclude that the metrics associated with $u_{0}$ and $\tilde{u}_{0}$, along with their potentials, agree up to natural conformal and diffeomorphism gauges; see Section 2.
The next step is to break these gauge invariances. Using unique continuation and the specific functional dependence of the potential on the metric $g$, we deduce (see Lemma 2.2 and (2.15)) that
(1.6) $|\nabla u_{0}|=|\nabla\tilde{u}_{0}|\circ\phi,$
where $\phi:\Omega\to\Omega$ is a diffeomorphism that fixes the boundary.
To strengthen this to full gradient equality, we employ the second linearization, which yields the integral identity:
(1.7) $0=\int_{\Omega}T\cdot\bigg($ $\nabla v_{0}\,g(\nabla v_{1},\nabla v_{2})$
$+\nabla v_{1}\left(\Delta_{g}(v_{0}v_{2})-v_{2}\Delta_{g}v_{0}-v_{0}\Delta_{g}v_{2}\right)$
$+\nabla v_{2}\left(\Delta_{g}(v_{1}v_{0})-v_{0}\Delta_{g}v_{1}-v_{1}\Delta_{g}v_{0}\right)\bigg)dV_{g},$
where $T$ is a
$T:=\left(\frac{\lambda J_{\phi}\phi^{*}\nabla\tilde{u}_{0}}{1+\lvert\nabla\tilde{u}_{0}\rvert^{2}|_{\phi}}-\frac{\nabla u_{0}}{1+\lvert\nabla u_{0}\rvert^{2}}\right)\lvert g\rvert^{-1/2},$
and $v_{k}$ ($k=0,1,2$) are solutions to the first linearized equation. The analysis of this identity relies on carefully constructed CGO solutions [GT11b], where we choose two solutions without critical points and one with a critical point. Using asymptotic analysis techniques similar to [CLLO24, LW23], we ultimately conclude that (see (4.4))
(1.8) $\nabla u_{0}=\phi^{*}\nabla\tilde{u}_{0}.$
The final step is to show that (1.6) and (1.8) together imply that $\phi$ is the identity map on $\Omega$. This follows from a uniqueness theorem for conformal mappings, which forces $\phi=\text{id}$, see Section 4.2. Consequently, $\nabla u_{0}=\nabla\tilde{u}_{0}$ throughout $\Omega$, and thus $H=\tilde{H}$ as required.
2. First linearization
To make notation more convenient, we let $F\colon\mathbb{R}^{2}\to\mathbb{R}^{2},$
$F(p):=\frac{p}{(1+\lvert p\rvert^{2})^{1/2}}$
and thus (1.1) can be written as
(2.1) $\nabla\cdot F(\nabla u)=H\text{ in }\Omega.$
We assume that there exists a solution $u_{0}\in C^{2,\alpha}(\Omega)$ to (2.1) corresponding to boundary value $f_{0}\in C^{2,\alpha}(\partial\Omega)$ and then we linearize at this solution. Let $f_{1},f_{2}\in C^{\infty}(\partial\Omega)$, $\epsilon=(\epsilon_{1},\epsilon_{2})$ and $\epsilon_{1},\epsilon_{2}>0,$ be such that $f=f_{0}+\epsilon_{1}f_{1}+\epsilon_{2}f_{2}$ is sufficiently close to $f_{0}$. Now the first linearization, denoted by
$v_{1}=\partial_{\epsilon_{1}}u|_{\epsilon=0},$
satisfies the first linearized equation
(2.2) $\partial_{a}(\partial_{b}F^{a}(\nabla u)|_{u=u_{0}}\partial_{\epsilon}\partial_{b}u|_{\epsilon=0})=\partial_{a}(\partial_{b}F^{a}(\nabla u_{0})\partial_{b}v)=0\text{ in }\Omega,$
where
(2.3) $\partial_{b}F^{a}(p)=\frac{\delta^{ab}}{(1+\lvert p\rvert^{2})^{1/2}}-\frac{p_{a}p_{b}}{(1+\lvert p\rvert^{2})^{3/2}}.$
From now on, we choose to put the indexes up and define the Riemannian metric
(2.4) $g^{ab}=\frac{1}{(1+\lvert\nabla u_{0}\rvert^{2})^{1/2}}\left(\delta^{ab}-\frac{\partial^{a}u_{0}\partial^{b}u_{0}}{1+\lvert\nabla u_{0}\rvert^{2}}\right)=\partial_{b}F^{a}(\nabla u_{0}).$
For a proof that $g^{ab}$ is indeed positive definite symmetric matrix we refer to [GT01, Chapter 10]. Thus the first linearization is the elliptic equation
(2.5) $\partial_{a}(g^{ab}\partial_{b}v)=0\text{ in }\Omega.$
For boundary values $f$ close to $f_{0}$, the boundary value problem
$\left\{\begin{array}[]{ll}\nabla\cdot F(\nabla u)=H,&\text{in}\,\,\Omega\\ u=f_{0}+f,&\text{on}\,\,\partial\Omega\end{array}\right$.
is well-posed and the DN map is smooth in the Fréchet sense:
Proposition 2.1.
Let $n\geq 2$, $\Omega\subset\mathbb{R}^{n}$ be a bounded domain with smooth boundary, $H\in C^{\infty}(\overline{\Omega})$ and $f_{0}\in C^{\infty}(\partial\Omega)$. Assume that $u_{0}\in C^{\infty}(\overline{\Omega})$ is a solution to
$\left\{\begin{array}[]{ll}\nabla\cdot F(\nabla u_{0})=H,&\text{in}\,\,\Omega\\ u_{0}=f_{0},&\text{on}\,\,\partial\Omega.\end{array}\right$.
Then there are $\delta,C>0$ such that for any
$f\in B_{\delta}=\{f\in C^{\infty}(\overline{\Omega}):\lVert f_{0}-f\rVert_{C^{\infty}(\overline{\Omega})}\leq\delta\}$
there exists a unique small solution $u\in C^{\infty}(\overline{\Omega})$ of
$\left\{\begin{array}[]{ll}\nabla\cdot F(\nabla u)=H,&\text{in}\,\,\Omega\\ u=f_{0}+f,&\text{on}\,\,\partial\Omega\end{array}\right$.
so that $\lVert u_{0}-u\rVert_{C^{\infty}(\overline{\Omega})}\leq C\delta$. Furthermore there are $C^{\infty}$ maps, in the Fréchet sense,
$S$ $\colon B_{\delta}\to C^{\infty}(\overline{\Omega}),\quad f\mapsto u$
$\Lambda_{H}$ $\colon B_{\delta}\to C^{\infty}(\overline{\Omega}),\quad f\mapsto\partial_{\nu}u|_{\partial\Omega}.$
The proof of this is similar to the one in [LL24] and we omit the proof. We note that the linearized equation (2.5) is well-posed (since it does not contain a zeroth order term) and thus we do not need to make assumptions regarding the well-posedness of the linearization of (1.1). If $H$ is small enough compared to the isoperimetric constant of $\Omega$ and the mean curvature of the boundary $\partial\Omega$, then $u_{0}$ exists and the solution $u$ corresponding to boundary value $f+f_{0}$ is globally unique. For these facts, see [GT01, Theorem 16.10, and (16.61)].
We modify (2.2) to a Laplace-Beltrami type equation. For $\gamma=\lvert g\rvert^{-1/2}$ we have
$0=\nabla\cdot\left(g^{-1}\nabla(\gamma^{-1/2}v)\right)=\gamma^{-1/2}(\Delta_{g}+q)v$
where $q=\dfrac{\Delta_{g}(\gamma^{1/2})}{\gamma^{1/2}}$. Thus $v$ is a solution of the boundary value problem
(2.6) $\left\{\begin{array}[]{ll}\nabla\cdot\left(g^{-1}\nabla v\right)=0,&\text{in}\,\,\Omega\\ v=f,&\text{on}\,\,\partial\Omega\end{array}\right$.
if and only if $\hat{v}=\gamma^{1/2}v$ is a solution to
(2.7) $\left\{\begin{array}[]{ll}(\Delta_{g}+q)\hat{v}=0,&\text{in}\,\,\Omega\\ \hat{v}=\gamma^{1/2}f,&\text{on}\,\,\partial\Omega.\end{array}\right$.
[…]
(2.8) $g=\lambda\phi^{*}\tilde{g}\quad\text{and that}\quad q=\lambda^{-1}\phi^{*}\tilde{q}$
with $\lambda|_{\partial\Omega}=1$ and $\phi|_{\partial\Omega}=\operatorname{Id}$, see [CLLO24].
Let us then denote by $v$ and $\tilde{v}$ the solutions to the first linearizations $(\Delta_{g}+q)v=0$ and $(\Delta_{\tilde{g}}+\tilde{q})\tilde{v}=0$ corresponding to a given boundary value $f\in C^{2,\alpha}(\partial\Omega)$. Now
(2.9) $(\Delta_{g}+q)(\tilde{v}\circ\phi)=(\Delta_{\lambda\phi^{*}\tilde{g}}+\lambda^{-1}\phi^{*}\tilde{q})(\tilde{v}\circ\phi)$
$=\lambda^{-1}(\Delta_{\phi^{*}\tilde{g}}+\phi^{*}\tilde{q})(\tilde{v}\circ\phi)=\lambda^{-1}\phi^{*}(\Delta_{\tilde{g}}\tilde{v}+\tilde{q}\tilde{v})=0.$
Since $\phi|_{\partial\Omega}=\mathrm{Id}$, we have $v=\tilde{v}\circ\phi$ by the uniqueness of solutions to the Dirichlet problem.
Using the Sylvester determinant rule and (2.4) we have the following
$\det(\lambda^{-1}(D\phi)^{-1}\tilde{g}^{-1}(D\phi)^{-T})$ $=\lambda^{-2}J_{\phi}^{-2}\frac{1}{1+\lvert\nabla\tilde{u}_{0}\rvert^{2}|_{\phi}}\left(1-\frac{\lvert\nabla\tilde{u}_{0}\rvert^{2}|_{\phi}}{1+\lvert\nabla\tilde{u}_{0}\rvert^{2}|_{\phi}}\right)$
$=\lambda^{-2}J_{\phi}^{-2}\frac{1}{(1+\lvert\nabla\tilde{u}_{0}\rvert^{2}|_{\phi})^{2}}.$
On the other hand
$\det(g^{-1})$ $=\frac{1}{1+\lvert\nabla u_{0}\rvert^{2}}\left(1-\frac{\lvert\nabla u_{0}\rvert^{2}}{1+\lvert\nabla u_{0}\rvert^{2}}\right)$
$=\frac{1}{(1+\lvert\nabla u_{0}\rvert^{2})^{2}}.$
Thus, the identity
$g^{-1}=\lambda^{-1}(D\phi)^{-1}\tilde{g}^{-1}|_{\phi}(D\phi)^{-T}$
gives
$(\lambda J_{\phi})^{-2}=\frac{(1+\lvert\nabla\tilde{u}_{0}\rvert^{2}|_{\phi})^{2}}{(1+\lvert\nabla u_{0}\rvert^{2})^{2}},$
or equivalently
(2.10) $\lambda J_{\phi}=\frac{1+\lvert\nabla u_{0}\rvert^{2}}{1+\lvert\nabla\tilde{u}_{0}\rvert^{2}|_{\phi}}.$
For the next result we recall that
[…]
where
[…]
Lemma 2.2.
Let $\Omega\subset\mathbb{R}^{2}$ be a domain with smooth boundary, $g,\tilde{g}$ be Riemannian metrics on $\Omega$, $\phi\colon\Omega\to\Omega$ a diffeomorphism such that $\phi|_{\partial\Omega}=\operatorname{Id}$ and $\lambda\colon\Omega\to(0,\infty)$ such that $\lambda|_{\partial\Omega}=1$. Assume that $g=\lambda\phi^{*}\tilde{g}$,
(2.12) $\det(g^{-1}\tilde{g}|_{\phi})=1\quad\text{to first order on }\partial\Omega$
and
(2.13) $q=\lambda^{-1}\phi^{*}\tilde{q}$,
where $q$ and $\tilde{q}$ are as in (2.11). Then
(2.14) $J_{\phi}\lambda=1.$
Proof.
By using the conformal invariance of the Laplacian, $\Delta_{cg}v=c^{-1}\Delta_{g}v$, together with (2.13) and $g=\lambda\phi^{*}\tilde{g}$ we have that
$q=\lvert g\rvert^{1/4}\Delta_{g}(\lvert g\rvert^{-1/4})$ $=\lambda^{-1}\phi^{*}\tilde{q}=\lambda^{-1}\phi^{*}\left(\lvert\tilde{g}\rvert^{1/4}\Delta_{\tilde{g}}(\lvert\tilde{g}\rvert^{-1/4})\right)$
$=\lambda^{-1}\lvert\tilde{g}\rvert^{1/4}|_{\phi}\Delta_{\phi^{*}\tilde{g}}(\lvert\tilde{g}\rvert^{-1/4}|_{\phi})$
$=\lambda^{-1}\lvert\tilde{g}\rvert^{1/4}|_{\phi}\Delta_{\lambda^{-1}g}(\lvert\tilde{g}\rvert^{-1/4}|_{\phi})$
$=\lvert\tilde{g}\rvert^{1/4}|_{\phi}\Delta_{g}(\lvert\tilde{g}\rvert^{-1/4}|_{\phi}).$
Let us denote $D=\lvert g\rvert^{-1/4}$ and $\tilde{D}=\lvert\tilde{g}\rvert^{1/4}|_{\phi}$. Then the previous equality becomes
[…]
This implies that
$\lvert\Delta_{g}(D-\tilde{D})\rvert=\lvert\tilde{D}\Delta_{g}D\rvert\lvert\frac{D-\tilde{D}}{D\tilde{D}}\rvert\leq C\lvert D-\tilde{D}\rvert$
for some positive constant $C$. By (2.12), we have that $D$ and $\tilde{D}$ agree to first order on $\partial\Omega$. It then follows by the standard unique continuation for the Laplace equation (see e.g [KSU11, Theorem B.1.]) that
D=D~ $\tilde{D}\ \text{ or equivalently }\ \lvert g\rvert=\lvert\tilde{g}\rvert|_{\phi}$.
By the assumption $g=\lambda\phi^{*}\tilde{g}$ and $\lvert g\rvert=\lvert\tilde{g}\rvert|_{\phi}$, we obtain
$\lvert g\rvert=\lvert\lambda\phi^{*}\tilde{g}\rvert=\lambda^{2}J_{\phi}^{2}\lvert\tilde{g}\rvert|_{\phi}=\lambda^{2}J_{\phi}^{2}\lvert g\rvert$
which then implies (2.14). ∎
Now from (2.14) and (2.10) we get
(2.15) $\lvert\nabla u_{0}\rvert=\lvert\nabla\tilde{u}_{0}\rvert|_{\phi}.$
3. Boundary information
Before moving to the second linearization in this section we discuss the needed information on the boundary. As seen in the previous section, the metric $g$ introduced in (2.4) depends on the initial solution $u_{0}$. Thus in order to have knowledge of $g$ on the boundary, we need to have knowledge of $u_{0}$ on the boundary.
First of all, since we know $u_{0}$ and $\partial_{\nu}u_{0}$ on the boundary, we know $\nabla u_{0}$ on the boundary. Thus we know $g$ on the boundary to zeroth order.
In Section 6 we will need that the conformal diffeomorphism $\phi$ from the previous section is the identity to high order on the boundary which follows from [CLLT23, Lemma 5.1]. We only need to check that $\phi$ satisfies the properties required by that Lemma. From (2.9) we know that $\phi$ is a morphism of solutions. We also know that $\Lambda_{H}=\Lambda_{\tilde{H}}$ and combining this with the fact that $g,\tilde{g}$ agree on the boundary further implies that the DN maps for the first linearizations agree.
Thus in order to use [CLLT23, Lemma 5.1] we need to have that $g,\tilde{g}$ agree on the boundary in all orders. This follows, if we can show that $u_{0}$ and $\tilde{u}_{0}$ agree on the boundary in all orders.
We know $H|_{\partial\Omega}=\tilde{H}|_{\partial\Omega}$ and thus
(3.1) $\nabla\cdot\left(\frac{\nabla u_{0}}{(1+\lvert\nabla u_{0}\rvert^{2})^{1/2}}\right)\bigg|_{\partial\Omega}=\nabla\cdot\left(\frac{\nabla\tilde{u}_{0}}{(1+\lvert\nabla\tilde{u}_{0}\rvert^{2})^{1/2}}\right)\bigg|_{\partial\Omega}.$
Now
$\nabla\cdot\left(\frac{\nabla u_{0}}{(1+\lvert\nabla u_{0}\rvert^{2})^{1/2}}\right)=\frac{\Delta u_{0}}{(1+\lvert\nabla u_{0}\rvert^{2})^{1/2}}-\frac{\nabla^{2}u_{0}(\nabla u_{0},\nabla u_{0})}{(1+\lvert\nabla u_{0}\rvert^{2})^{3/2}}$
which can be divided into terms containing tangential derivatives of $u_{0}$, which we already know on $\partial\Omega$, and terms containing only normal derivatives. For the moment let us assume that the normal direction is in the direction $x_{2}$. The ones with only normal derivatives are
$\frac{\partial^{2}_{2}u_{0}}{(1+\lvert\nabla u_{0}\rvert^{2})^{1/2}}-\frac{\partial^{2}_{2}u_{0}\partial_{2}u_{0}\partial_{2}u_{0}}{(1+\lvert\nabla u_{0}\rvert^{2})^{3/2}}=\partial^{2}_{2}u_{0}\left(\frac{1}{(1+\lvert\nabla u_{0}\rvert^{2})^{1/2}}-\frac{(\partial_{2}u_{0})^{2}}{(1+\lvert\nabla u_{0}\rvert^{2})^{3/2}}\right).$
Then from (3.1) it follows
$\frac{\partial^{2}_{2}u_{0}}{(1+\lvert\nabla u_{0}\rvert^{2})^{1/2}}\left(1-\frac{(\partial_{2}u_{0})^{2}}{1+\lvert\nabla u_{0}\rvert^{2}}\right)=\frac{\partial^{2}_{2}\tilde{u}_{0}}{(1+\lvert\nabla\tilde{u}_{0}\rvert^{2})^{1/2}}\left(1-\frac{(\partial_{2}\tilde{u}_{0})^{2}}{1+\lvert\nabla\tilde{u}_{0}\rvert^{2}}\right)$
and using $\nabla u_{0}=\nabla\tilde{u}_{0}$ on $\partial\Omega$ implies
$\left(\partial^{2}_{2}u_{0}-\partial^{2}_{2}\tilde{u}_{0}\right)\left(1-\frac{(\partial_{2}u_{2})^{2}}{1+\lvert\nabla u_{0}\rvert^{2}}\right)=0.$
This is true when $\partial^{2}_{2}u_{0}=\partial^{2}_{2}\tilde{u}_{0}$ since otherwise we would have $(\partial_{2}u_{0})^{2}=-1$. Thus from $H|_{\partial\Omega}=\tilde{H}|_{\partial\Omega}$ it follows $\nabla^{2}u_{0}|_{\partial\Omega}=\nabla^{2}\tilde{u}_{0}|_{\partial\Omega}$.
For higher order derivatives of $u_{0}$ we again only need to focus on terms with only higher order normal derivatives. Since we know $\partial_{2}H|_{\partial\Omega}=\partial_{2}\tilde{H}|_{\partial\Omega}$, we compute $\partial_{2}H$:
$\partial_{2}H$ $=\frac{\partial_{2}\Delta u_{0}}{(1+\lvert\nabla u_{0}\rvert^{2})^{1/2}}$
$-(1+\lvert\nabla u_{0}\rvert^{2})^{-3/2}\left(\partial^{3}_{211}u_{0}(\partial_{1}u_{0})^{2}+\partial^{3}_{212}u_{0}\partial_{1}u_{0}\partial_{2}u_{0}+\partial^{2}_{221}u_{0}\partial_{1}u_{0}\partial_{2}u_{0}+\partial^{3}_{2}u_{0}(\partial_{2}u_{0})^{2}\right)$
$+R(\nabla^{2}u_{0},\nabla u_{0})$
where the term $R$ does not include third order derivatives. Again we are interested in the term with only $\partial_{2}$ derivatives in them:
$\partial^{3}_{2}u_{0}\left(\frac{1}{(1+\lvert\nabla u_{0}\rvert^{2})^{1/2}}-\frac{(\partial_{2}u_{0})^{2}}{(1+\lvert\nabla u_{0}\rvert^{2})^{3/2}}\right).$
This is similar to situation above and thus we would have $\partial^{3}_{2}u_{0}=\partial^{3}_{2}\tilde{u}_{0}$. Hence we get that the third order derivatives of $u_{0}$ and $\tilde{u}_{0}$ agree on $\partial\Omega$. Proceeding inductively we obtain that $u_{0}$ and $\tilde{u}_{0}$ agree to high order on $\partial\Omega$.
Thus we can use [CLLT23, Lemma 5.1] to say that $\phi$ is the identity on $\partial\Omega$ to infinite order. This together with (2.10) implies that $\lambda$ is equal to one on $\partial\Omega$ to infinite order.
4. Second linearization
Using the same notation as in Section 2, we calculate the second linearization of (1.1). Now the second linearization is
(4.1) $\partial_{a}\big(\partial_{p_{c}p_{b}}F^{a}(\nabla u_{0})\partial_{c}v_{1}\partial_{b}v_{2}+\partial_{p_{b}}F^{a}(\nabla u_{0})\partial_{b}w\big)=0,$
where $w=\partial_{\epsilon_{1}\epsilon_{2}}u|_{\epsilon=0}$. Next we calculate $\partial_{p_{c}p_{b}}F^{a}(p)$:
$\partial_{p_{c}p_{b}}F^{a}(p)$ $=-\frac{p_{c}}{(1+\lvert p\rvert^{2})^{3/2}}\left(\delta_{ab}-\frac{1}{1+\lvert p\rvert^{2}}p_{a}p_{b}\right)$
$+\frac{1}{(1+\lvert p\rvert^{2})^{1/2}}\left(\frac{2p_{a}p_{b}p_{c}}{(1+\lvert p\rvert^{2})^{2}}-\frac{\delta_{ac}p_{b}+\delta_{bc}p_{a}}{1+\lvert p\rvert^{2}}\right)$
$=-\frac{p_{c}}{(1+\lvert p\rvert^{2})^{3/2}}\left(\delta_{ab}-\frac{1}{1+\lvert p\rvert^{2}}p_{a}p_{b}\right)$
$+\frac{1}{(1+\lvert p\rvert^{2})^{3/2}}\left(-p_{b}\left(\delta_{ac}-\frac{p_{a}p_{c}}{1+\lvert p\rvert^{2}}\right)-p_{a}\left(\delta_{bc}-\frac{p_{b}p_{c}}{1+\lvert p\rvert^{2}}\right)\right).$
Using (2.3) and (2.4) gives
$\partial_{p_{c}p_{b}}F^{a}(\nabla u_{0})=-\frac{1}{1+\lvert\nabla u_{0}\rvert^{2}}(\partial_{a}u_{0}g^{bc}+\partial_{b}u_{0}g^{ac}+\partial_{c}u_{0}g^{ab}).$
Define, again with indices up,
$C^{abc}:=\partial_{p_{c}p_{b}}F^{a}(\nabla u_{0}).$
Notice that $\partial_{p_{c}p_{b}}F^{a}(p)$ is symmetric.
4.1. Integral identity
Here we derive an integral identity for the second linearization. Integrating (4.1) against a solution $v_{0}$ of the first linearization to obtain
$0=$ $\int_{\Omega}\partial_{a}\big(C^{abc}\partial_{c}v_{1}\partial_{b}v_{2}+g_{0}^{ab}\partial_{b}w\big)v_{0}\,dx$
$=$ $-\int_{\Omega}C^{abc}\partial_{c}v_{1}\partial_{b}v_{2}\partial_{a}v_{0}\,dx-\int_{\Omega}g_{0}^{ab}\partial_{b}w\partial_{a}v_{0}\,dx$
$+\int_{\partial\Omega}\big(C^{abc}\partial_{c}v_{1}\partial_{b}v_{2}+g_{0}^{ab}\partial_{b}w\big)v_{0}\nu_{a}\,dS$
$=$ $-\int_{\Omega}C^{abc}\partial_{c}v_{1}\partial_{b}v_{2}\partial_{a}v_{0}\,dx$
$+\int_{\partial\Omega}\big(C^{abc}\partial_{c}v_{1}\partial_{b}v_{2}+g_{0}^{ab}\partial_{b}w\big)v_{0}\nu_{a}-wg_{0}^{ab}\partial_{a}v_{0}\nu_{b}\,dS$
$=-I+B.$
For the moment let us only focus on II:
$I$ $=\int_{\Omega}-\frac{1}{1+\lvert\nabla u_{0}\rvert^{2}}(\partial_{a}u_{0}g^{bc}+\partial_{b}u_{0}g^{ac}+\partial_{c}u_{0}g^{ab})\partial_{c}v_{1}\partial_{b}v_{2}\partial_{a}v_{0}\,dx$
$=\int_{\Omega}-\frac{1}{1+\lvert\nabla u_{0}\rvert^{2}}(\partial_{a}u_{0}\partial_{a}v_{0}g(\nabla v_{1},\nabla v_{2})+\partial_{b}u_{0}\partial_{b}v_{2}g(\nabla v_{0},\nabla v_{1})+\partial_{c}u_{0}\partial_{c}v_{1}g(\nabla v_{0},\nabla v_{2}))\,dx$
$=\int_{\Omega}-\frac{\nabla u_{0}}{1+\lvert\nabla u_{0}\rvert^{2}}\cdot(\nabla v_{0}g(\nabla v_{1},\nabla v_{2})+\nabla v_{1}g(\nabla v_{0},\nabla v_{2})+\nabla v_{2}g(\nabla v_{0},\nabla v_{1}))\,dx.$
Next consider two sources $H,\tilde{H}$ as in the previous section. Now by making a change of coordinates with $\phi$
$\tilde{I}$ $=\int_{\Omega}-\frac{1}{1+\lvert\nabla\tilde{u}_{0}\rvert^{2}}(\partial_{a}\tilde{u}_{0}\tilde{g}^{bc}+\partial_{b}\tilde{u}_{0}\tilde{g}^{ac}+\partial_{c}\tilde{u}_{0}\tilde{g}^{ab})\partial_{c}\tilde{v}_{1}\partial_{b}\tilde{v}_{2}\partial_{a}\tilde{v}_{0}\,dx$
$=\int_{\Omega}-\frac{1}{1+\lvert\nabla\tilde{u}_{0}\rvert^{2}|_{\phi}}((\partial_{a}\tilde{u}_{0})|_{\phi}\tilde{g}^{bc}|_{\phi}+(\partial_{b}\tilde{u}_{0})|_{\phi}\tilde{g}^{ac}|_{\phi}+(\partial_{c}\tilde{u}_{0})|_{\phi}\tilde{g}^{ab}|_{\phi})$
$\times[D\phi^{-1}]^{c^{\prime}}_{c}[D\phi^{-1}]^{b^{\prime}}_{b}[D\phi^{-1}]^{a^{\prime}}_{a}\partial_{c^{\prime}}v_{1}\partial_{b^{\prime}}v_{2}\partial_{a^{\prime}}v_{0}J_{\phi}\,dx$
$=\int_{\Omega}-\frac{\lambda J_{\phi}}{1+\lvert\nabla\tilde{u}_{0}\rvert^{2}|_{\phi}}((\phi^{*}\nabla\tilde{u}_{0})^{a^{\prime}}g^{b^{\prime}c^{\prime}}+(\phi^{*}\nabla\tilde{u}_{0})^{b^{\prime}}g^{a^{\prime}c^{\prime}}+(\phi^{*}\nabla\tilde{u}_{0})^{c^{\prime}}g^{a^{\prime}b^{\prime}})$
$\times\partial_{c^{\prime}}v_{1}\partial_{b^{\prime}}v_{2}\partial_{a^{\prime}}v_{0}\,dx$
$=\int_{\Omega}-\frac{\lambda J_{\phi}\phi^{*}\nabla\tilde{u}_{0}}{1+\lvert\nabla\tilde{u}_{0}\rvert^{2}|_{\phi}}\cdot(\nabla v_{0}g(\nabla v_{1},\nabla v_{2})+\nabla v_{1}g(\nabla v_{0},\nabla v_{2})+\nabla v_{2}g(\nabla v_{0},\nabla v_{1}))\,dx,$
where we used (2.8) for $g^{-1},\tilde{g}^{-1}$ i.e. $g^{-1}=\lambda^{-1}(D\phi)^{-1}\tilde{g}^{-1}(D\phi)^{-T}$. Let
(4.2) $T:=\left(\frac{\lambda J_{\phi}\phi^{*}\nabla\tilde{u}_{0}}{1+\lvert\nabla\tilde{u}_{0}\rvert^{2}|_{\phi}}-\frac{\nabla u_{0}}{1+\lvert\nabla u_{0}\rvert^{2}}\right)\lvert g\rvert^{-1/2}.$
Then
(4.3) $I-\tilde{I}$ $=\int_{\Omega}T\cdot(\nabla v_{0}g(\nabla v_{1},\nabla v_{2})+\nabla v_{1}g(\nabla v_{0},\nabla v_{2})+\nabla v_{2}g(\nabla v_{0},\nabla v_{1}))\,dV_{g}$
$=\int_{\Omega}T\cdot\bigg(\nabla v_{0}g(\nabla v_{1},\nabla v_{2})+\nabla v_{1}\left(\Delta_{g}(v_{0}v_{2})-v_{2}\Delta_{g}v_{0}-v_{0}\Delta_{g}v_{2}\right)$
$+\nabla v_{2}\left(\Delta_{g}(v_{1}v_{0})-v_{0}\Delta_{g}v_{1}-v_{1}\Delta_{g}v_{0}\right)\bigg)\,dV_{g}.$
Now if we can show that $T=0$ which further implies, using (2.10) and (2.15),
(4.4) $\phi^{*}\nabla\tilde{u}_{0}=\nabla u_{0}.$
4.2. Showing that $\phi$ is the identity
From the first linearization we know that $\lambda D\phi^{-1}\tilde{g}^{-1}|_{\phi}D\phi^{-T}=g^{-1}$ or equivalently (using the explicit formula (2.4) for $g$ and $\tilde{g}$)
$\lambda D\phi^{-1}D\phi^{-T}-\frac{D\phi^{-1}(\nabla\tilde{u}_{0}\otimes\nabla\tilde{u}_{0})|_{\phi}D\phi^{-T}}{1+\lvert\nabla\tilde{u}_{0}\rvert^{2}|_{\phi}}=I-\frac{\nabla u_{0}\otimes\nabla u_{0}}{1+\lvert\nabla u_{0}\rvert^{2}}.$
Here we used (2.15). Using the above together with (4.4) and (2.15) implies
$\lambda D\phi^{-1}D\phi^{-T}=I\quad\text{or}\quad D\phi^{-1}D\phi^{-T}=\lambda^{-1}I.$
That is $\phi$ is a conformal mapping in Euclidean space. Combining this with $\phi|_{\partial\Omega}=\mathrm{Id}$ a Liouville typer result (see e.g. [Lio97, Proposition $3.$]) gives that $\phi$ is the identity in $\Omega$.
4.3. Boundary terms
When deriving the integral identity above we encountered the following integral on the boundary
$B=\int_{\partial\Omega}\big(C^{abc}\partial_{c}v_{1}\partial_{b}v_{2}+g_{0}^{ab}\partial_{b}w\big)v_{0}\nu_{a}-wg_{0}^{ab}\partial_{a}v_{0}\nu_{b}\,dS.$
We know the boundary values of $u_{0}$ and $\partial_{\nu}u_{0}$, hence we know $\nabla u_{0}$ on the boundary. Thus by the definition of $g$ we know $g$ on the boundary. This further implies that we know $C$ on the boundary. Since we know that the DN map is smooth and we know the DN map related to (1.1), we thus know the DN map of the linearized equation. This also uses that $g$ is known on the boundary.
Using these and that $\phi$ is the identity on the boundary to high order (this follows from [CLLO24, Lemma 6.1], see Section 3) we have that $\lambda$ is equal to 11 on the boundary. Putting all of the above together we can conclude that
$B-\tilde{B}=0.$
5. Complex geometrical optics solutions
In this section, we construct CGO solutions for the first linearized equation (2.6) $\nabla\cdot\left(g^{-1}\nabla v\right)=0$. As already noted in (2.7) it is sufficient to construct solutions to the Shrödinger equation
$(\Delta_{g}+q)\mathbf{v}=0\text{ in }\Omega$
[…]
Let us then recall the standard holomorphic calculus on Riemannian surfaces. The complexified cotangent bundle $\mathbb{C}T^{*}\widetilde{M}$ has the splitting
$\mathbb{C}T^{*}\widetilde{M}=T^{*}_{1,0}\widetilde{M}\oplus T^{*}_{0,1}\widetilde{M}$
determined by the eigenspaces of the Hodge star operator $\star$. In local holomorphic coordinate $z$ the space $T^{*}_{1,0}\widetilde{\Sigma}$ is spanned by $dz$ and $T^{*}_{0,1}\widetilde{\Sigma}$ is spanned by $d\overline{z}$. The invariant definitions of $\overline{\partial}$ and $\partial$ operators are $\pi_{0,1}d$ and $\pi_{1,0}d$ respectively. In holomorphic coordinates $z=(x,y)$ these operators are
[…]
By [GT11b, Proposition 2.1] there is a right inverse $\overline{\partial}^{-1}$ for $\overline{\partial}$ in the sense that
[…]
which is the Hermitean adjoint of $\overline{\partial}$. In holomorphic coordinates the operator $\overline{\partial}^{*}$ is just $\partial$. If $\Sigma$ admits global holomorphic coordinates, then in those coordinates $\overline{\partial}^{-1}$ and $\overline{\partial}^{*-1}$ equal
[…]
with the understanding that the function $\mathbf{u}\in L^{p}(\Sigma)$ in the coordinates are continued onto $\mathbb{C}$ in $L^{p}$ (and similarly if $\mathbf{u}$ has higher regularity). We also define
$\overline{\partial}_{\psi}^{-1}:=\mathcal{R}\overline{\partial}^{-1}e^{-2i\psi/h}\mathcal{E}\quad\text{and}\quad\overline{\partial}_{\psi}^{*-1}:=\mathcal{R}\overline{\partial}^{*-1}e^{2i\psi/h}\mathcal{E},$
[…]
(5.1) $\begin{split}\lVert\overline{\partial}_{\psi}^{-1}f\rVert_{L^{q}(M)}&\leq Ch^{1/q}\lVert f\rVert_{W^{1,p}(M,T_{0,1}^{*}M)}\end{split}$
Moreover, there is $\epsilon>0$ such that
(5.2) $\begin{split}\lVert\overline{\partial}_{\psi}^{-1}f\rVert_{L^{2}(M)}&\leq Ch^{1/2+\epsilon}\lVert f\rVert_{W^{1,p}(M,T_{0,1}^{*}M)}\end{split}$
We have similar estimates for $\overline{\partial}_{\psi}^{*-1}f$.
The CGOs on a Riemannian surface $(\Sigma,g)$ for the equation $(\Delta_{g}+q)\mathbf{v}=0$ have the form
(5.3) $\mathbf{v}=e^{\Phi/h}(a+r_{h}).$
Here $\Phi$ is holomorphic function, which is Morse (i.e. Hessian at possible critical points of $\Phi$ is invertible). Here also $a$ is a holomorphic function and $r=r_{h}$ is a correction term. The precise form of $r_{h}$ is
(5.4) $r_{h}=-\overline{\partial}_{\psi}^{-1}\sum_{j=0}^{\infty}T_{h}^{j}\overline{\partial}_{\psi}^{*-1}(qa),$
where the operators $\overline{\partial}_{\psi}^{-1}$ and $\overline{\partial}_{\psi}^{*-1}$ were defined above and $T_{h}$ is defined as
(5.5) $T_{h}:=-\overline{\partial}_{\psi}^{*-1}q\overline{\partial}_{\psi}^{-1}.$
We also denote
(5.6) $s_{h}=\sum_{j=0}^{\infty}T_{h}^{j}\overline{\partial}_{\psi}^{*-1}(qa)$
so that $r_{h}=-\overline{\partial}_{\psi}^{-1}s_{h}$.
By taking complex conjugate and changing the sign of the phase function, we also have CGO solutions with antiholomorphic phase of the form
(5.7) $\mathbf{v}=e^{-\overline{\Phi}/h}(b+\tilde{r}_{h}),$
where $\overline{\Phi}$ is antiholomorphic Morse function, $b$ antiholomorphic and $\tilde{r}_{h}$ given by:
(5.8) $\tilde{r}_{h}=-\partial_{\psi}^{-1}\sum_{j=0}^{\infty}\tilde{T}_{h}^{j}(\partial_{\psi}^{*-1}(qa))=-\partial_{\psi}^{-1}\tilde{s}_{h}.$
Let us then recall estimates from [GT11b, GT11a, CLLO24, LW23]. If $\Phi$ has a critical point, then
[…]
for any $p\in[2,\infty)$ and $0<\epsilon\leq\epsilon_{p}$ depending on $p$. Moreover, we have
(5.10) $\lVert T_{h}\rVert_{L^{r}\to L^{r}}=O(h^{1/r})\text{ and }\lVert T_{h}\rVert_{L^{2}\to L^{2}}=O(h^{1/2-\epsilon}),$
for any $0<\epsilon<1/2$ and $r>2$. See [CLLO24, Section 4.1] for the above. If $f\in C^{\infty}(\Sigma)$ vanishes to order $1$ on $\partial\Sigma$ and $\hat{\psi}$ is the imaginary part of holomorphic Morse function on $\Sigma$, then
(5.11) $\int_{\Sigma}e^{i\hat{\psi}/h}fr_{h}=-\int_{\Sigma}e^{i\hat{\psi}/h}f\overline{\partial}_{\psi}^{-1}s_{h}=\int_{\Sigma}\overline{\partial}^{-1}(e^{i\hat{\psi}/h}f)e^{-2i\psi/h}s_{h}=O(h^{1+\epsilon})$
[…]
If $\Phi$ has no critical points, we have the better estimates
(5.12) $\|r_{h}\|_{p}+\|s_{h}\|_{p}+\|dr_{h}\|_{p}\leq Ch$
for all $p\in(1,\infty)$. Also, in this case we have the expansion
(5.13) $\overline{\partial}_{\psi}^{-1}f=e^{-2i\psi/h}\frac{ih}{2}\frac{f}{\bar{\partial}\psi}+\frac{ih}{2}\overline{\partial}^{-1}\left(e^{-2i\psi/h}\bar{\partial}\left(\frac{f}{\bar{\partial}\psi}\right)\right),$
[…]
(5.14) $H\nabla^{2}r_{h}||_{L^{2}}=O(1)$
see [CLLO24, Section 4.1] and [LW23, Eq. (3.14)].
We have similar estimates for $\tilde{r}_{h}$, $\tilde{s}_{h}$ and $\tilde{T}_{h}$, and expansion for $\overline{\partial}_{\psi}^{*-1}$, to the ones above.
6. Proving that $T=0$
Here we assume that $T$ vanishes to high order on the boundary $\partial\Omega$. The integral identity (4.3) holds for arbitrary solutions of the first linearization (2.6). We will construct these to be of the form $v=\gamma^{-1/2}\hat{v}$, $\gamma=\lvert g\rvert^{-1/2}$, where $\hat{v}$ is a solution of (2.7). Thus we can modify by noting that for these solutions
$\Delta_{g}v_{k}$ $=\Delta_{g}(\gamma^{-1/2}\hat{v}_{k})=\Delta_{g}(\gamma^{-1/2})\hat{v}_{k}+2g\left(\nabla(\gamma^{-1/2}),\nabla\hat{v}_{k}\right)+\gamma^{-1/2}\Delta_{g}\hat{v}_{k}$
$=\Delta_{g}(\gamma^{-1/2})\hat{v}_{k}+2g\left(\nabla(\gamma^{-1/2}),\nabla\hat{v}_{k}\right)-\gamma^{-1/2}q\hat{v}_{k}$
$=2g\left(\nabla(\gamma^{-1/2}),\nabla\hat{v}_{k}\right)+\left(\Delta_{g}(\gamma^{-1/2})-\gamma^{-1/2}q\right)\hat{v}_{k},$
where $q=\lvert g\rvert^{1/4}\Delta_{g}(\lvert g\rvert^{-1/4})$:
$I-\tilde{I}$ $=\sum_{k=1}^{2}\int_{\Omega}T^{k}\Big(\partial_{k}v_{0}g(\nabla v_{1},\nabla v_{2})$
$+\partial_{k}v_{1}(\Delta_{g}(v_{0}v_{2})-\gamma^{-1/2}\hat{v}_{2}\Delta_{g}(\gamma^{-1/2}\hat{v}_{0})-\gamma^{-1/2}\hat{v}_{0}\Delta_{g}(\gamma^{-1/2}\hat{v}_{2}))$
$+\partial_{k}v_{2}(\Delta_{g}(v_{0}v_{1})-\gamma^{-1/2}\hat{v}_{0}\Delta_{g}(\gamma^{-1/2}\hat{v}_{1})-\gamma^{-1/2}\hat{v}_{1}\Delta_{g}(\gamma^{-1/2}\hat{v}_{0}))\Big)\,dV_{g}$
$=\sum_{k=1}^{2}\int_{\Omega}T^{k}\partial_{k}v_{0}g(\nabla v_{1},\nabla v_{2})$
$+T^{k}\partial_{k}v_{1}\left(\Delta_{g}(v_{0}v_{2})+2\left(g\left(\nabla(\gamma^{-1/2}),\nabla\hat{v}_{0}\right)\hat{v}_{2}+g\left(\nabla(\gamma^{-1/2}),\nabla\hat{v}_{2}\right)\hat{v}_{0}\right)+2Q\hat{v}_{0}\hat{v}_{2}\right)$
$+T^{k}\partial_{k}v_{2}\left(\Delta_{g}(v_{0}v_{1})+2\left(g\left(\nabla(\gamma^{-1/2}),\nabla\hat{v}_{0}\right)\hat{v}_{1}+g\left(\nabla(\gamma^{-1/2}),\nabla\hat{v}_{1}\right)\hat{v}_{0}\right)+2Q\hat{v}_{0}\hat{v}_{1}\right)\,dV_{g}$
$:=I_{1}+I_{2}+I_{3}.$
Here $Q:=\Delta_{g}(\gamma^{-1/2})-\gamma^{-1/2}q$. We choose the functions $\hat{v}_{k}$ as follows: Let $\Psi,\Phi,a_{k}$ be holomorphic Morse functions, $k=1,2$, and $a_{0}$ antiholomorphic such that $\pm\Psi+\Phi$ has no critical points and $\Phi$ has critical points, see Section 5. Then
$\hat{v}_{0}$ $=e^{-2\overline{\Phi}/h}(a_{0}+r_{0})$
$\hat{v}_{1}$ $=e^{(\Psi+\Phi)/h}(a_{1}+r_{1})$
$\hat{v}_{2}$ $=e^{(-\Psi+\Phi)/h}(a_{2}+r_{2})$
and
$v_{k}=\gamma^{-1/2}\hat{v}_{k}\quad\text{ for }k=0,1,2.$
At this point it is beneficial to use holomorphic coordinates around an interior point $p\in\Omega$. In these coordinates $z=x+iy$, we have $g(z)=c(z)dzd\overline{z}$ where $c$ is a smooth function, $dz=dx+idy$ and $d\overline{z}=dx-idy$. Recall also the operators
(6.1) $\partial=\frac{1}{2}(\partial_{x}-i\partial_{y})\quad\text{and}\quad\overline{\partial}=\frac{1}{2}(\partial_{x}+i\partial_{y}),$
thus $\nabla_{(x,y)}=(\partial+\overline{\partial},i(\partial-\overline{\partial}))$. In these coordinates
(6.2) $g(\nabla u,\nabla v)=2c^{-1}(\partial u\overline{\partial}v+\overline{\partial}u\partial v),\quad\Delta_{g}u=-2ic^{-1}\overline{\partial}\partial u\quad\text{and}\quad dV_{g}=c(z)\frac{i}{2}dzd\overline{z}.$
Since $v_{1}$ and $v_{2}$ have a holomorphic phase, they act similarly in the asymptotic analysis, and thus it is enough to do analysis for $I_{2}$ as the analysis for $I_{3}$ is the same and in the end we get the same terms from stationary phase (see Lemmas 6.1, 6.2, 6.3 below). Using the above, we have for $I_{2}$ after integrating by parts
$I_{2}$ $=\sum_{k=1}^{2}\int_{\Omega}T^{k}\partial_{k}v_{1}\overline{\partial}\partial(v_{0}v_{2})\,dzd\overline{z}$
$+i\int_{\Omega}cT^{k}\partial_{k}v_{1}\left(g\left(\nabla(\gamma^{-1/2}),\nabla\hat{v}_{0}\right)\hat{v}_{2}+g\left(\nabla(\gamma^{-1/2}),\nabla\hat{v}_{2}\right)\hat{v}_{0}\right)\,dzd\overline{z}$
$+i\int_{\Omega}cT^{k}\partial_{k}v_{1}Q\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=I_{2,1}+I_{2,2}+I_{2,3}$
and similarly for $I_{3}$. Here we make a slight abuse of notation by writing again $T^{k}$ for its representation in holomorphic coordinates.
In the analysis of the above integrals we will use stationary phase arguments for terms including $T$. For this we need that $T$ vanishes to high order on the boundary $\partial\Omega$. Recall that
$T:=\left(\frac{\lambda J_{\phi}\phi^{*}\nabla\tilde{u}_{0}}{1+\lvert\nabla\tilde{u}_{0}\rvert^{2}|_{\phi}}-\frac{\nabla u_{0}}{1+\lvert\nabla u_{0}\rvert^{2}}\right)\lvert g\rvert^{-1/2}.$
As explained in Section 3 we have that $\Phi$ is the identity, $\lambda=1$ and $u_{0}=\tilde{u}_{0}$ to high order on $\partial\Omega$. Thus $T$ vanishes to high order on $\partial\Omega$. We are now ready to prove the following:
Lemma 6.1.
Define $A_{5}=2T^{k}c^{-1}\overline{\partial}(\gamma^{-1/2})+\overline{\partial}T^{k}\gamma^{-1/2}$. Then
$\lim_{h\to 0}hI_{2,1}=\sum_{k=1}^{2}Ci^{k-1}(A_{5}\gamma^{-1})|_{z=z_{0}}$
for some constant $C>0$.
Proof.
Integrating by parts in $I_{2,1}$ we have
$I_{2,1}$ $=\sum_{k=1}^{2}\int_{\Omega}\overline{\partial}\partial(T^{k}\partial_{k}v_{1})v_{0}v_{2}\,dzd\overline{z}$
$=\sum_{k=1}^{2}\int_{\Omega}\overline{\partial}\partial(T^{k})\partial_{k}v_{1}v_{0}v_{2}+T^{k}\overline{\partial}\partial(\partial_{k}v_{1})v_{0}v_{2}+\overline{\partial}T^{k}\partial\partial_{k}(v_{1})v_{0}v_{2}+\partial T^{k}\overline{\partial}\partial_{k}v_{1}v_{0}v_{2}\,dzd\overline{z}$
$=\sum_{k=1}^{2}\int_{\Omega}\overline{\partial}\partial(T^{k})\partial_{k}v_{1}v_{0}v_{2}+T^{k}\partial_{k}(2g\left(\nabla(\gamma^{-1/2}),\nabla\hat{v}_{1}\right)+Q\hat{v}_{1})v_{0}v_{2}$
$+\overline{\partial}T^{k}\partial\partial_{k}(v_{1})v_{0}v_{2}+\partial T^{k}\overline{\partial}\partial_{k}v_{1}v_{0}v_{2}\,dzd\overline{z}.$
Using (6.1) and (6.2) we can group the integral $I_{2,1}$ as follows:
$I_{2,1}$ $=\sum_{k=1}^{2}\int_{\Omega}\Big(\overline{\partial}\partial(T^{k})\partial_{k}(\gamma^{-1/2})+T^{k}\partial_{k}Q+\overline{\partial}T^{k}\partial\partial_{k}(\gamma^{-1/2})+\partial T^{k}\overline{\partial}\partial_{k}(\gamma^{-1/2})\Big)\hat{v}_{1}v_{0}v_{2}$
$+\Big(i^{k-1}\overline{\partial}\partial(T^{k})\gamma^{-1/2}+2T^{k}\partial_{k}(c^{-1})\overline{\partial}(\gamma^{-1/2})+2T^{k}c^{-1}\partial_{k}\overline{\partial}(\gamma^{-1/2})+i^{k-1}T^{k}Q$
$+\overline{\partial}T^{k}\partial_{k}(\gamma^{-1/2})+i^{k-1}\overline{\partial}T^{k}\partial(\gamma^{-1/2})+i^{k-1}\partial T^{k}\overline{\partial}(\gamma^{-1/2})\Big)\partial\hat{v}_{1}v_{0}v_{2}$
$+\Big((-i)^{k-1}\overline{\partial}\partial(T^{k})\gamma^{-1/2}+2T^{k}\partial_{k}(c^{-1})\partial(\gamma^{-1/2})+2T^{k}c^{-1}\partial_{k}\partial(\gamma^{-1/2})+(-i)^{k-1}T^{k}Q$
$+(-i)^{k-1}\overline{\partial}T^{k}\partial(\gamma^{-1/2})+\partial T^{k}\partial_{k}(\gamma^{-1/2})+(-i)^{k-1}\partial T^{k}\overline{\partial}(\gamma^{-1/2})\Big)\overline{\partial}\hat{v}_{1}v_{0}v_{2}$
$+\Big(2T^{k}c^{-1}\partial(\gamma^{-1/2})+\partial T^{k}\gamma^{-1/2}\Big)\overline{\partial}\partial_{k}\hat{v}_{1}v_{0}v_{2}$
$+\Big(2T^{k}c^{-1}\overline{\partial}(\gamma^{-1/2})+\overline{\partial}T^{k}\gamma^{-1/2}\Big)\partial\partial_{k}\hat{v}_{1}v_{0}v_{2}\,dzd\overline{z}$
$=\sum_{k=1}^{2}\int_{\Omega}A_{1}\hat{v}_{1}v_{0}v_{2}+A_{2}\partial\hat{v}_{1}v_{0}v_{2}+A_{3}\overline{\partial}\hat{v}_{1}v_{0}v_{2}+A_{4}\overline{\partial}\partial_{k}\hat{v}_{1}v_{0}v_{2}+A_{5}\partial\partial_{k}\hat{v}_{1}v_{0}v_{2}\,dzd\overline{z}.$
All of the terms $A_{j}$ depend on the summation index $k$.
In the following we do not write the summation over $k$. Next we further analyze $I_{2,1}$ term by term, starting from the term including $A_{1}$:
$\int_{\Omega}A_{1}\hat{v}_{1}v_{0}v_{2}\,dzd\overline{z}$ $=\int_{\Omega}\gamma^{-1}A_{1}e^{(2\Phi-2\Phi)/h}(a_{0}+r_{0})(a_{1}+r_{1})(a_{2}+r_{2})\,dzd\overline{z}=O(1)$
as $h\to 0$, since the functions $\gamma,A_{1},a_{j}$ are bounded in $\Omega$ and by the estimates for $r_{j}$ (see (5.9) and (5.12)). Next the term with $A_{2}$:
$\int_{\Omega}A_{2}\partial\hat{v}_{1}v_{0}v_{2}$
$=\frac{1}{h}\int_{\Omega}\gamma^{-1}A_{2}\partial(\Psi+\Phi)e^{(2\Phi-2\overline{\Phi})/h}(a_{0}+r_{0})(a_{1}+r_{1})(a_{2}+r_{2})\,dzd\overline{z}$
$+\int_{\Omega}\gamma^{-1}A_{2}e^{(\Psi+\Phi)/h}(\partial a_{1}+\partial r_{1})\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=\frac{1}{h}\int_{\Omega}\gamma^{-1}A_{2}\partial(\Psi+\Phi)e^{(2\Phi-2\overline{\Phi})/h}a_{0}a_{1}a_{2}\,dzd\overline{z}+O(h^{-1/2+\epsilon})+O(1),$
where the term $O(h^{-1/2+\epsilon})$ comes from the fact that all the terms include at least one $r_{j}$ and the term $O(1)$ similarly as for the term with $A_{1}$. The last term will be $O(1)$ by stationary phase and hence
(6.3) $\int_{\Omega}A_{2}\partial\hat{v}_{1}v_{0}v_{2}=O(h^{-1/2+\epsilon}).$
The term with $A_{3}$ is a little bit simpler since $\overline{\partial}\Phi=\overline{\partial}\Psi=\overline{\partial}a_{1}=0$ by the holomorphicity of these functions. Thus
$\int_{\Omega}A_{3}\overline{\partial}\hat{v}_{1}v_{0}v_{2}\,dzd\overline{z}=\int_{\Omega}\gamma^{-1}A_{3}e^{(\Psi+\Phi)/h}\overline{\partial}r_{1}\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}=O(h)$
because of the term $\overline{\partial}r_{1}$, see (5.12). Continuing with $A_{4}$:
$\int_{\Omega}A_{4}\overline{\partial}\partial_{k}\hat{v}_{1}v_{0}v_{2}\,dzd\overline{z}$ $=\int_{\Omega}\gamma^{-1}A_{4}\left(i^{k-1}\overline{\partial}\partial\hat{v}_{1}\hat{v}_{0}\hat{v}_{2}+(-i)^{k-1}\overline{\partial}^{2}\hat{v}_{1}\hat{v}_{0}\hat{v}_{2}\right)\,dzd\overline{z}$
$=\int_{\Omega}\gamma^{-1}A_{4}\frac{i^{k-2}}{2}cq\hat{v}_{1}\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$+\int_{\Omega}\gamma^{-1}A_{4}(-i)^{k-1}\overline{\partial}^{2}\left(e^{(\Psi+\Phi)/h}(a_{1}+r_{1})\right)\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=O(1)+\int_{\Omega}\gamma^{-1}A_{4}(-i)^{k-1}\overline{\partial}^{2}\left(e^{(\Psi+\Phi)/h}(a_{1}+r_{1})\right)\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
and this can be seen by the same arguments as for the term with $A_{1}$. For the rest, we notice that each time the operator $\overline{\partial}$ hits $e^{(\Psi+\Phi)/h}$ we get zero since $\overline{\partial}\Phi=\overline{\partial}\Psi=0$. Also $\overline{\partial}a_{1}=0$, hence
$\int_{\Omega}\gamma^{-1}A_{4}(-i)^{k-1}\overline{\partial}^{2}\left(e^{(\Psi+\Phi)/h}(a_{1}+r_{1})\right)\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=\int_{\Omega}\gamma^{-1}A_{4}(-i)^{k-1}e^{(\Psi+\Phi)/h}\overline{\partial}^{2}r_{1}\hat{v}_{0}\hat{v}_{2}$
$=O(1)$
by the same arguments as before and using the Calderón-Zygmund type estimate in (5.14) for $\overline{\partial}^{2}r_{1}$.
The term with $A_{5}$ is the most complicated of these. First of all
$\int_{\Omega}A_{5}\partial\partial_{k}\hat{v}_{1}v_{0}v_{2}\,dzd\overline{z}$
$=\int_{\Omega}A_{5}i^{k-1}\partial^{2}\hat{v}_{1}v_{0}v_{2}\,dzd\overline{z}+\int_{\Omega}A_{5}(-i)^{k-1}\partial\overline{\partial}\hat{v}_{1}v_{0}v_{2}\,dzd\overline{z}$
$=\int_{\Omega}A_{5}\gamma^{-1}i^{k-1}\partial\left(\frac{1}{h}\partial(\Psi+\Phi)e^{(\Psi+\Phi)/h}(a_{1}+r_{1})+e^{(\Psi+\Phi)/h}(\partial a_{1}+\partial r_{1})\right)\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$+O(1)$
by the same arguments as for the integral with $A_{4}$. Expanding further
$\int_{\Omega}A_{5}\partial\partial_{k}\hat{v}_{1}v_{0}v_{2}\,dzd\overline{z}$
$=\int_{\Omega}A_{5}\gamma^{-1}i^{k-1}\frac{1}{h^{2}}(\partial(\Psi+\Phi))^{2}e^{(\Psi+\Phi)/h}(a_{1}+r_{1})\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$+\int_{\Omega}A_{5}\gamma^{-1}i^{k-1}\frac{1}{h}\partial^{2}(\Psi+\Phi)e^{(\Psi+\Phi)/h}(a_{1}+r_{1})\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$+2\int_{\Omega}A_{5}\gamma^{-1}i^{k-1}\frac{1}{h}\partial(\Psi+\Phi)e^{(\Psi+\Phi)/h}(\partial a_{1}+\partial r_{1})\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$+\int_{\Omega}A_{5}\gamma^{-1}i^{k-1}e^{(\Psi+\Phi)/h}(\partial^{2}a_{1}+\partial^{2}r_{1})\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}+O(1).$
Focusing first on the terms with $\frac{1}{h}$, we get that they are $O(h^{-1/2+\epsilon})$ by stationary phase (for terms with no remainders) and that the remainders $r_{k}$ satisfy the estimates (5.9), (5.12). For the ones with $\frac{1}{h^{2}}$ we have three kinds of terms: one without any $r_{k}$, ones with just one $r_{k}$ and ones with two or more $r_{k}$:
$\int_{\Omega}A_{5}\gamma^{-1}i^{k-1}\frac{1}{h^{2}}(\partial(\Psi+\Phi))^{2}e^{(\Psi+\Phi)/h}(a_{1}+r_{1})\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=\frac{1}{h^{2}}\int_{\Omega}A_{5}\gamma^{-1}i^{k-1}(\partial(\Psi+\Phi))^{2}e^{(2\Phi-2\overline{\Phi})/h}a_{0}a_{1}a_{2}\,dzd\overline{z}$
$+\frac{1}{h^{2}}\int_{\Omega}A_{5}\gamma^{-1}i^{k-1}(\partial(\Psi+\Phi))^{2}e^{(2\Phi-2\overline{\Phi})/h}\Big(a_{1}a_{0}r_{2}+a_{1}r_{0}a_{2}+a_{0}r_{1}a_{2}$
$+a_{1}r_{0}r_{2}+a_{0}r_{1}r_{2}+r_{0}r_{1}a_{2}+r_{0}r_{1}r_{2}\Big)\,dzd\overline{z}.$
Because $\Psi$ and $\Phi$ were chosen so that $\partial(\Psi+\Phi)(z_{0})\neq 0$ we can use stationary phase to conclude that the first term is
$\frac{1}{h^{2}}\int_{\Omega}A_{5}\gamma^{-1}i^{k-1}(\partial(\Psi+\Phi))^{2}e^{(2\Phi-2\overline{\Phi})/h}a_{0}a_{1}a_{2}\,dzd\overline{z}=C\frac{1}{h}i^{k-1}(A_{5}\gamma^{-1})|_{z=z_{0}}+O(1)$
for some constant $C>0$. For the terms with at least two $r_{k}$, we can use the estimates (5.9), (5.12) for these to conclude that they are $O(h^{-1/2+\epsilon})$ as $h\to 0$. We are then left with the terms that have only one $r_{k}$. Let us consider the case with $r_{0}$. The other two are exactly the same and the estimates for $r_{1},r_{2}$ are better than for $r_{0}$. Now by the definition of $r_{0}$ we have
(6.4) $\frac{1}{h^{2}}\int_{\Omega}A_{5}\gamma^{-1}i^{k-1}(\partial(\Psi+\Phi))^{2}e^{(2\Phi-2\overline{\Phi})/h}a_{1}a_{2}r_{0}\,dzd\overline{z}$
$=-\frac{1}{h^{2}}\int_{\Omega}A_{5}\gamma^{-1}i^{k-1}(\partial(\Psi+\Phi))^{2}e^{(2\Phi-2\overline{\Phi})/h}a_{1}a_{2}\partial^{-1}_{\operatorname{Im}(-2\overline{\Phi})}s_{0}\,dzd\overline{z}$
$=O(h^{-1+\epsilon}).$
Here we used that one can integrate by parts with the operator $\partial^{-1}_{\psi}$, see (5.11). All in all we have
(6.5) $\int_{\Omega}A_{5}\partial\partial_{k}\hat{v}_{1}v_{0}v_{2}\,dzd\overline{z}=C\frac{1}{h}i^{k-1}(A_{5}\gamma^{-1})|_{z=z_{0}}+O(h^{-1+\epsilon})$
which concludes the proof. ∎
Now we move to analyze the integral $I_{2,2}$ (i.e. to (6.6)).
Lemma 6.2.
Define $B_{7}=i^{k-1}T^{k}\overline{\partial}(\gamma^{-1/2})$. Then
$\lim_{h\to 0}hI_{2,2}=\sum_{k=1}^{2}iC(B_{7}\gamma^{-1/2})|_{z=z_{0}}$
for some constant $C>0$.
Proof.
For $I_{2,2}$ we have, using the same calculations as for $I_{2,1}$,
(6.6) $-\frac{i}{2}I_{2,2}$ $=\sum_{k=1}^{2}\int_{\Omega}i^{k-1}T^{k}\partial(\gamma^{-1/2})\partial v_{1}\overline{\partial}\hat{v}_{0}\hat{v}_{2}+(-i)^{k-1}T^{k}\partial(\gamma^{-1/2})\overline{\partial}v_{1}\overline{\partial}\hat{v}_{0}\hat{v}_{2}$
$+i^{k-1}T^{k}\overline{\partial}(\gamma^{-1/2})\partial v_{1}\partial\hat{v}_{0}\hat{v}_{2}+(-i)^{k-1}T^{k}\overline{\partial}(\gamma^{-1/2})\overline{\partial}v_{1}\partial\hat{v}_{0}\hat{v}_{2}$
$+i^{k-1}T^{k}\partial(\gamma^{-1/2})\partial v_{1}\overline{\partial}\hat{v}_{2}\hat{v}_{0}+(-i)^{k-1}T^{k}\partial(\gamma^{-1/2})\overline{\partial}v_{1}\overline{\partial}\hat{v}_{2}\hat{v}_{0}$
$+i^{k-1}T^{k}\overline{\partial}(\gamma^{-1/2})\partial v_{1}\partial\hat{v}_{2}\hat{v}_{0}+(-i)^{k-1}T^{k}\overline{\partial}(\gamma^{-1/2})\overline{\partial}v_{1}\partial\hat{v}_{2}\hat{v}_{0}\,dzd\overline{z}$
$=\sum_{k=1}^{2}\int_{\Omega}B_{1}\partial v_{1}\overline{\partial}\hat{v}_{0}\hat{v}_{2}+B_{2}\overline{\partial}v_{1}\overline{\partial}\hat{v}_{0}\hat{v}_{2}+B_{3}\partial v_{1}\partial\hat{v}_{0}\hat{v}_{2}+B_{4}\overline{\partial}v_{1}\partial\hat{v}_{0}\hat{v}_{2}$
$+B_{5}\partial v_{1}\overline{\partial}\hat{v}_{2}\hat{v}_{0}+B_{6}\overline{\partial}v_{1}\overline{\partial}\hat{v}_{2}\hat{v}_{0}+B_{7}\partial v_{1}\partial\hat{v}_{2}\hat{v}_{0}+B_{8}\overline{\partial}v_{1}\partial\hat{v}_{2}\hat{v}_{0}\,dzd\overline{z}.$
Again the terms $B_{j}$ depend on $k$. The first term is
$\int_{\Omega}B_{1}\partial v_{1}\overline{\partial}\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=\int_{\Omega}B_{1}\partial(\gamma^{-1/2})\hat{v}_{1}\overline{\partial}\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}+\int_{\Omega}B_{1}\gamma^{-1/2}\partial\hat{v}_{1}\overline{\partial}\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=\int_{\Omega}B_{1}\partial(\gamma^{-1/2})e^{(2\Phi-2\overline{\Phi})/h}(a_{1}+r_{1})\Big(\frac{-2}{h}\overline{\partial}\overline{\Phi}(a_{0}+r_{0})+\overline{\partial}a_{0}+\overline{\partial}r_{0}\Big)(a_{2}+r_{2})\,dzd\overline{z}$
$+\int_{\Omega}B_{1}\gamma^{-1/2}e^{(2\Phi-2\overline{\Phi})/h}\Big(\frac{1}{h}\partial(\Psi+\Phi)(a_{1}+r_{1})+\partial a_{1}+\partial r_{1}\Big)$
$\times\Big(\frac{-2}{h}\overline{\partial}\overline{\Phi}(a_{0}+r_{0})+\overline{\partial}a_{0}+\overline{\partial}r_{0}\Big)(a_{2}+r_{2})\,dzd\overline{z}$
$=O(h^{-1/2+\epsilon})+O(1)$
by stationary phase and the estimates for $r_{k}$ (see (5.9), (5.12)). For the second term in (6.6) we have
$\int_{\Omega}B_{2}\overline{\partial}v_{1}\overline{\partial}\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=\int_{\Omega}B_{2}\overline{\partial}(\gamma^{-1/2})\hat{v}_{1}\overline{\partial}\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}+\int_{\Omega}B_{2}\gamma^{-1/2}\overline{\partial}\hat{v}_{1}\overline{\partial}\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=O(h^{-1/2}+\epsilon)$
$+\int_{\Omega}B_{2}\gamma^{-1/2}e^{(2\Phi-2\overline{\Phi})/h}\overline{\partial}r_{1}\Big(\frac{-2}{h}\overline{\partial}\overline{\Phi}(a_{0}+r_{0})+\overline{\partial}a_{0}+\overline{\partial}r_{0}\Big)(a_{2}+r_{2})\,dzd\overline{z}$
$=O(h^{-1/2+\epsilon})+O(1)$
where the first term is as above with $B_{1}$ and the second is handled with the estimate (5.12) for $\overline{\partial}r_{1}$. Next the term with $B_{3}$:
$\int_{\Omega}B_{3}\partial v_{1}\partial\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=\int_{\Omega}B_{3}\partial(\gamma^{-1/2})\hat{v}_{1}\partial\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}+\int_{\Omega}B_{3}\gamma^{-1/2}\partial\hat{v}_{1}\partial\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=\int_{\Omega}B_{3}\partial(\gamma^{-1/2})e^{(2\Phi-2\overline{\Phi})/h}\partial r_{0}(a_{1}+r_{1})(a_{2}+r_{2})\,dzd\overline{z}$
$+\int_{\Omega}B_{3}\gamma^{-1/2}e^{(2\Phi-2\overline{\Phi})/h}\Big(\frac{1}{h}\partial(\Psi+\Phi)(a_{1}+r_{1})+\partial a_{1}+\partial r_{1}\Big)\partial r_{0}(a_{2}+r_{2})\,dzd\overline{z}$
$=O(h^{1/2+\epsilon})+O(h^{-1/2+\epsilon})$
since all the terms have $\partial r_{0}$ in them (see (5.9)). For the next one we have
$\int_{\Omega}B_{4}\overline{\partial}v_{1}\partial\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=\int_{\Omega}B_{4}\overline{\partial}(\gamma^{-1/2})\hat{v}_{1}\partial\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}+\int_{\Omega}B_{4}\gamma^{-1/2}\overline{\partial}\hat{v}_{1}\partial\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=O(h^{1/2+\epsilon})+\int_{\Omega}B_{4}\gamma^{-1/2}e^{(2\Phi-2\overline{\Phi})/h}\overline{\partial}r_{1}\partial r_{0}(a_{2}+r_{2})\,dzd\overline{z}$
$=O(h^{1/2+\epsilon})+O(h^{3/2+\epsilon})$
similarly as above.
For the terms with $B_{5}$ to $B_{8}$ we only consider the term with $B_{7}$ in detail. For $B_{5},B_{6}$ they both have the term $\overline{\partial}\hat{v}_{2}=e^{(-\Psi+\Phi)/h}\overline{\partial}r_{2}$ and thus are $O(h^{\alpha})$ with $\alpha>-1/2$ (see (5.12)). For $B_{8}$ we have terms like $\hat{v}_{1}\partial\hat{v}_{2}\hat{v}_{0}$ and $\overline{\partial}\hat{v}_{1}\partial\hat{v}_{2}\hat{v}_{0}$. The first is similar to $A_{2}$ (see (6.3)) and the second one falls to the same category as $B_{5},B_{6}$. For $B_{7}$ we have
$\int_{\Omega}B_{7}\partial v_{1}\partial\hat{v}_{2}\hat{v}_{0}\,dzd\overline{z}$
$=\int_{\Omega}B_{7}\partial(\gamma^{-1/2})\hat{v}_{1}\partial\hat{v}_{2}\hat{v}_{0}\,dzd\overline{z}+\int_{\Omega}B_{7}\gamma^{-1/2}\partial\hat{v}_{1}\partial\hat{v}_{2}\,dzd\overline{z}$
$=O(h^{-1/2+\epsilon})$
$+\int_{\Omega}B_{7}\gamma^{-1/2}e^{(2\Phi-2\overline{\Phi})/h}\Big(\frac{1}{h}\partial(\Psi+\Phi)(a_{1}+r_{1})+\partial a_{1}+\partial r_{1}\Big)$
$\times\Big(\frac{1}{h}\partial(-\Psi+\Phi)(a_{2}+r_{2})+\partial a_{2}+\partial r_{2}\Big)(a_{0}+r_{0})\,dzd\overline{z},$
where the first integral was similar to (6.3). The rest we divide into three cases by the powers of $h$. If there is $h^{0}$, then using stationary phase or the estimates for $r_{k}$ (see (5.9), (5.12)), these terms are $O(1)$. Similarly for the terms with $h^{-1}$ and for the terms with $h^{-2}$ with at least two $r_{k}$ we get $O(h^{-1/2+\epsilon})$. Then we are left with the following:
$\int_{\Omega}B_{7}\partial v_{1}\partial\hat{v}_{2}\hat{v}_{0}\,dzd\overline{z}$
$=\frac{1}{h^{2}}\int_{\Omega}B_{7}\gamma^{-1/2}e^{(2\Phi-2\overline{\Phi})/h}\partial(\Psi+\Phi)\partial(-\Psi+\Phi)\big(a_{0}a_{1}a_{2}+a_{1}a_{2}r_{0}+a_{0}a_{2}r_{1}\big)\,dzd\overline{z}$
$+O(h^{-1/2+\epsilon}).$
For the last two terms we use the definitions of $r_{0}$ and $r_{1}$, as we did in (6.4), and for the first one we use stationary phase to obtain for $C>0$
(6.7) $\int_{\Omega}B_{7}\partial v_{1}\partial\hat{v}_{2}\hat{v}_{0}\,dzd\overline{z}=C\frac{1}{h}(B_{7}\gamma^{-1/2})|_{z=z_{0}}+O(h^{-1+\epsilon}).$
This ends the proof. ∎
Next we consider the last term in $I_{2}$.
Lemma 6.3.
For $I_{2,3}$ we have $\lim_{h\to 0}hI_{2,3}=0.$
Proof.
For $I_{2,3}$ we get the following:
$-iI_{2,3}$ $=\sum_{k=1}^{2}\int_{\Omega}i^{k-1}cT^{k}Q\partial v_{1}\hat{v}_{0}\hat{v}_{2}+(-i)^{k-1}cT^{k}Q\overline{\partial}v_{1}\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}$
$=\sum_{k=1}^{2}\int_{\Omega}C_{1}\partial v_{1}\hat{v}_{0}\hat{v}_{2}+C_{2}\overline{\partial}v_{1}\hat{v}_{0}\hat{v}_{2}\,dzd\overline{z}.$
These are at most $O(h^{-1/2+\epsilon})$ by the same arguments as for the terms with $A_{1},A_{2}$ and $A_{3}$ (see before and after equation (6.3)). ∎
Then we return to the integral $I_{1}$, that is to
$I_{1}=\sum_{k=1}^{2}\int_{\Omega}T^{k}\partial_{k}v_{0}g(\nabla v_{1},\nabla v_{2})\,dV_{g}.$
Lemma 6.4.
For $I_{1}$ we have $\lim_{h\to 0}hI_{1}=0.$
Proof.
We immediately use (6.2) to have
$\sum_{k=1}^{2}\int_{\Omega}T^{k}\partial_{k}v_{0}g(\nabla v_{1},\nabla v_{2})\,dV_{g}$
$=\sum_{k=1}^{2}i\int_{\Omega}T^{k}\left(i^{k-1}\partial v_{0}+(-i)^{k-1}\overline{\partial}v_{0}\right)\left(\partial v_{1}\overline{\partial}v_{2}+\overline{\partial}v_{1}\partial v_{2}\right)\,dzd\overline{z}$
$=\sum_{k=1}^{2}i\int_{\Omega}D_{1}\partial v_{0}\partial v_{1}\overline{\partial}v_{2}+D_{2}\partial v_{0}\overline{\partial}v_{1}\partial v_{2}+D_{3}\overline{\partial}v_{0}\partial v_{1}\overline{\partial}v_{2}+D_{4}\overline{\partial}v_{0}\overline{\partial}v_{1}\partial v_{2}\,dzd\overline{z}.$
Here all $D_{j}$ again depend on the summation index $k$.
Starting with the $D_{1}$ term we expand further (and as before we do not write the summation):
(6.8) $\int_{\Omega}D_{1}\partial v_{0}\partial v_{1}\overline{\partial}v_{2}\,dzd\overline{z}$
$=\int_{\Omega}\tilde{D}_{1}\hat{v}_{0}\hat{v}_{1}\hat{v}_{2}+\tilde{D}_{1}\hat{v}_{0}\hat{v}_{1}\overline{\partial}\hat{v}_{2}+\tilde{D}_{1}\hat{v}_{0}\partial\hat{v}_{1}\hat{v}_{2}+\tilde{D}_{1}\hat{v}_{0}\partial\hat{v}_{1}\overline{\partial}\hat{v}_{2}+\tilde{D}_{1}\partial\hat{v}_{0}\hat{v}_{1}\overline{\partial}\hat{v}_{2}$
$+\tilde{D}_{1}\partial\hat{v}_{0}\partial\hat{v}_{1}\hat{v}_{2}+\tilde{D}_{1}\partial\hat{v}_{0}\hat{v}_{1}\hat{v}_{2}+\tilde{D}_{1}\partial\hat{v}_{0}\partial\hat{v}_{1}\overline{\partial}\hat{v}_{2}\,dzd\overline{z}.$
Here we have abused notation using $\tilde{D}_{1}$ for all the terms that do not depend on $h$. The first six terms are similar to $A_{1},A_{3},A_{2},B_{5},B_{4}$ and $B_{3}$, in that order, and they are $O(h^{-1/2+\epsilon})$. For the first remaining term we have
$\int_{\Omega}\tilde{D}_{1}\partial\hat{v}_{0}\hat{v}_{1}\hat{v}_{2}\,dzd\overline{z}$
$=\int_{\Omega}\tilde{D}_{1}\partial\hat{v}_{0}\hat{v}_{1}\hat{v}_{2}\,dzd\overline{z}$
$=-\frac{2}{h}\int_{\Omega}\tilde{D}_{1}\partial\overline{\Phi}e^{-2\overline{\Phi}/h}(a_{0}+r_{0})\hat{v}_{1}\hat{v}_{2}+\tilde{D}_{1}e^{(2\Phi-2\overline{\Phi})/h}\partial r_{0}(a_{1}+r_{1})(a_{2}+r_{2})\,dzd\overline{z}.$
The first term is $O(1)$ by stationary phase and the fact that $\partial\overline{\Phi}(z_{0})=0$ and the second term is $O(h^{1/2+\epsilon})$ because of the term $\partial r_{0}$ (see (5.9)). For the last term in (6.8) we use that $\partial a_{0}=\overline{\partial}a_{2}=\overline{\partial}(-\Psi+\Phi)=0=\overline{\partial}\overline{\Phi}(z_{0})$ and stationary phase to have
$\int_{\Omega}\tilde{D}_{1}\partial\hat{v}_{0}\partial\hat{v}_{1}\overline{\partial}\hat{v}_{2}\,dzd\overline{z}$
$=\int_{\Omega}e^{(2\Phi-2\overline{\Phi})/h}\partial r_{0}\left(\frac{1}{h}\partial(\Psi+\Phi)(a_{1}+r_{1})+\partial a_{1}+\partial r_{1}\right)\partial r_{2}\,dzd\overline{z}$
$=O(h^{1/2+\epsilon}).$
In the last equality we used the estimates for $\partial r_{0},\partial r_{2}$ (see (5.9), (5.12)).
The term with $D_{2}$ goes similarly as above with the term $D_{1}$ because $v_{1}$ and $v_{2}$ have holomorphic phases.
Moving to the term with $D_{3}$ we have (using similar notation as with $D_{1}$)
$\int_{\Omega}D_{3}\overline{\partial}v_{0}\partial v_{1}\overline{\partial}v_{2}\,dzd\overline{z}$
$=\int_{\Omega}\tilde{D}_{3}\hat{v}_{0}\hat{v}_{1}\hat{v}_{2}+\tilde{D}_{3}\hat{v}_{0}\hat{v}_{1}\overline{\partial}\hat{v}_{2}+\tilde{D}_{3}\hat{v}_{0}\partial\hat{v}_{1}\hat{v}_{2}+\tilde{D}_{3}\hat{v}_{0}\partial\hat{v}_{1}\overline{\partial}\hat{v}_{2}+\tilde{D}_{3}\overline{\partial}\hat{v}_{0}\hat{v}_{1}\overline{\partial}\hat{v}_{2}$
$+\tilde{D}_{3}\overline{\partial}\hat{v}_{0}\hat{v}_{1}\hat{v}_{2}+\tilde{D}_{3}\overline{\partial}\hat{v}_{0}\partial\hat{v}_{1}\overline{\partial}\hat{v}_{2}\,dzd\overline{z}.$
Again the first six terms are similar to ones before. These are $O(h^{-1/2+\epsilon})$ by the arguments made for $A_{1},A_{3},A_{2},B_{5},B_{6}$ and $B_{1}$. The first remaining term is
$\int_{\Omega}\tilde{D}_{3}\overline{\partial}\hat{v}_{0}\hat{v}_{1}\hat{v}_{2}\,dzd\overline{z}$
$=\int_{\Omega}e^{(2\Phi-2\overline{\Phi})/h}\left(\left(-\frac{2}{h}\overline{\partial}\overline{\Phi}(a_{0}+r_{0})+\overline{\partial}a_{0}+\overline{\partial}r_{0}\right)(a_{1}+r_{1})(a_{2}+r_{2})\right)\,dzd\overline{z}$
$=O(1)$
by stationary phase and the estimates for $r_{k}$ (see (5.9), (5.12)). The last remaining term is, by using again that the functions $a_{2},\Phi$ and $\Psi$ are holomorphic,
$\int_{\Omega}\tilde{D}_{3}\overline{\partial}\hat{v}_{0}\partial\hat{v}_{1}\overline{\partial}\hat{v}_{2}\,dzd\overline{z}$
$=\int_{\Omega}\tilde{D}_{3}e^{(2\Phi-2\overline{\Phi})/h}\left(-\frac{2}{h}\overline{\partial}\overline{\Phi}(a_{0}+r_{0})+\overline{\partial}a_{0}+\overline{\partial}r_{0}\right)$
$\times\left(\frac{1}{h}\partial(\Psi+\Phi)(a_{1}+r_{1})+\partial a_{1}+\partial r_{1}\right)\overline{\partial}r_{2}\,dzd\overline{z}$
$=-\frac{2}{h^{2}}\int_{\Omega}\tilde{D}_{3}e^{(2\Phi-2\overline{\Phi})/h}\overline{\partial}\overline{\Phi}\partial(\Psi+\Phi)a_{1}a_{0}\overline{\partial}r_{2}\,dzd\overline{z}+O(h^{-1/2+\epsilon})$
because of the estimates for $r_{1},r_{0}$ and $\overline{\partial}r_{2}$ (see (5.9), (5.12)). For the last term, we recall that (see (5.4) - (5.6))
$r_{2}$ $=-\overline{\partial}^{-1}_{\operatorname{Im}(-\Psi+\Phi)}s_{2}=\overline{\partial}^{-1}e^{i\operatorname{Im}(-\Psi+\Phi)/h}s_{2}$
$s_{2}$ $=-\overline{\partial}_{\operatorname{Im}(-\Psi+\Phi)}^{*-1}(qa)-\sum_{j=1}^{\infty}T_{2}^{j}\overline{\partial}_{\operatorname{Im}(-\Psi+\Phi)}^{*-1}(qa)$
and thus in holomorphic coordinates, denoting $\eta=\operatorname{Im}(-\Psi+\Phi)$,
$\overline{\partial}r_{2}$ $=-e^{i\eta/h}s_{2}$
$=e^{i\eta/h}\partial^{-1}(e^{i\eta/h}qa)+e^{i\eta/h}\sum_{j=1}^{\infty}T_{2}^{j}\partial^{-1}(e^{i\eta/h}qa)$
$=e^{i\eta/h}\left(\frac{ih}{2}e^{i\eta/h}\frac{qa}{\partial\eta}+\frac{ih}{2}\partial^{-1}\left(e^{i\eta/h}\partial\left(\frac{qa}{\partial\eta}\right)\right)\right)+e^{i\eta/h}\sum_{j=1}^{\infty}T_{2}^{j}\partial^{-1}(e^{\eta/h}qa).$
Here we used the expansion (5.13). Now
$-\frac{2}{h^{2}}\int_{\Omega}\tilde{D}_{3}e^{(2\Phi-2\overline{\Phi})/h}\overline{\partial}\overline{\Phi}\partial(\Psi+\Phi)a_{1}a_{0}\frac{ih}{2}e^{2i\eta/h}\frac{qa}{\partial\eta}\,dzd\overline{z}=O(1)$
by stationary phase. Recall that
$\partial^{-1}\left(e^{i\eta/h}\partial\left(\frac{qa}{\partial\eta}\right)\right)=O(h^{1/2+\epsilon}),\quad\partial^{-1}(e^{i\eta/h}qa)=O(h)$
(by the estimate (5.2) and the expansion (5.13)) and $T_{h}=O(h^{1/2-\epsilon})$. Then the remaining terms in $\int_{\Omega}\tilde{D}_{3}\overline{\partial}\hat{v}_{0}\partial\hat{v}_{1}\overline{\partial}\hat{v}_{2}\,dzd\overline{z}$ are $O(h^{3/2-\epsilon})$
Again the term with $D_{4}$ goes similarly as above with the term $D_{3}$ because $v_{1}$ and $v_{2}$ have holomorphic phases. ∎
Recall the definition of $T$ in (4.2) and that the arguments in Lemmas 6.1, 6.2, 6.3 can be applied to the integral $I_{3}$ with the same results. Then we have the following for $T$:
Lemma 6.5.
The components $T^{1},T^{2}$ of $T$ satisfy the elliptic equation
(6.9) $\partial\overline{\partial}(T^{1}+iT^{2})+E\partial(T^{1}+iT^{2})+\partial E(T^{1}+iT^{2})=0\quad\text{in }\Omega$
for a smooth function $E$.
Proof.
Multiplying $I-\tilde{I}$ by $h$ and taking the limit as $h\to 0$ we get after using Lemmas 6.1, 6.2, 6.3, 6.4
$\sum_{k=1}^{2}2Ci^{k-1}(A_{5}\gamma^{-1})|_{z=z_{0}}+i\tilde{C}(B_{7}\gamma^{-1/2})|_{z=z_{0}}=0.$
Recall that $A_{5}=2T^{k}c^{-1}\overline{\partial}(\gamma^{-1/2})+\overline{\partial}T^{k}\gamma^{-1/2}$ and $B_{7}=i^{k-1}T^{k}\overline{\partial}(\gamma^{-1/2})$. Doing the above argument for a dense set of points implies that $T_{1},T_{2}$ satisfy
$\sum_{k=1}^{2}i^{k-1}\overline{\partial}T^{k}+\frac{i^{k-1}}{2C}\gamma^{3/2}\left(4Cc^{-1}\overline{\partial}(\gamma^{-1/2})\gamma^{-1}+i\tilde{C}\overline{\partial}(\gamma^{-1/2})\gamma^{-1/2}\right)T^{k}=0.$
Letting $E:=\frac{1}{2C}\gamma^{3/2}\left(4Cc^{-1}\overline{\partial}(\gamma^{-1/2})\gamma^{-1}+i\tilde{C}\overline{\partial}(\gamma^{-1/2})\gamma^{-1/2}\right)$ we have
$\overline{\partial}(T^{1}+iT^{2})+E(T^{1}+iT^{2})=0.$
Operating with $\overline{\partial}$ on both sides gives (6.9). ∎
Next we choose solutions $v_{k}=\gamma^{-1/2}\hat{v}_{k}$, $k=0,1,2$, such that
$\hat{v}_{0}$ $=e^{2\Phi/h}(\overline{a}_{0}+r_{0})$
$\hat{v}_{1}$ $=e^{-(\overline{\Psi}+\overline{\Phi})/h}(\overline{a}_{1}+r_{1})$
$\hat{v}_{2}$ $=e^{-(-\overline{\Psi}+\overline{\Phi})/h}(\overline{a}_{2}+r_{2})$
where $\Phi,\Psi,a_{k}$, $k=0,1,2$, are as before. Then we look at the integral $I-\tilde{I}$ with these solutions. The analysis is in most parts the same. The differences come from the terms $I_{2,1}$ and $I_{2,2}$ (and of course $I_{3,1}$ and $I_{3,2}$ but they are identical to $I_{2,1}$ and $I_{2,2}$). With the solutions above we get
$\lim_{h\to 0}hI_{2,1}$ $=\sum_{k=1}^{2}C(-i)^{k-1}(A_{4}\gamma^{-1})|_{z=z_{0}}$
$\lim_{h\to 0}hI_{2,2}$ $=\sum_{k=1}^{2}i\tilde{C}(B_{6}\gamma^{-1/2})|_{z=z_{0}},$
where
$A_{4}=\Big(2T^{k}c^{-1}\overline{\partial}(\gamma^{-1/2})+\overline{\partial}T^{k}\gamma^{-1/2}\Big),\quad B_{6}=(-i)^{k-1}T^{k}\partial(\gamma^{-1/2}).$
Then, similarly as for Lemma 6.5 we have the following:
Lemma 6.6.
The components $T^{1},T^{2}$ of $T$ satisfy the elliptic equation
$\partial\overline{\partial}(T^{1}-iT^{2})+\tilde{E}\partial(T^{1}-iT^{2})+\partial\tilde{E}(T^{1}-iT^{2})=0\quad\text{in }\Omega$
for a smooth function $\tilde{E}$.
Since we assume that $T$ is known on the boundary to high order, we get by combining Lemma 6.5 and Lemma 6.6 with unique continuation that $T^{1}=T^{2}=0$.
Acknowledgment. The authors would like to thank Niko Jokela for helpful discussion. T.L. was supported by the Academy of Finland (Centre of Excellence in Inverse Modeling and Imaging, grant numbers 284715 and 309963). J.N. was supported by the Finnish Centre of Excellence in Inverse Modelling and Imaging (Academy of Finland Grant 284715), the Research Council of Finland (Flagship of Advanced Mathematics for Sensing Imaging and Modelling grants 359208 and 359183) and by the Emil Aaltonen Foundation.

### v2

@@
 Abstract.
-We extend the study of inverse problems for minimal surfaces by considering the inverse source problem for the prescribed mean curvature equation
+We extend the recent study of inverse problems for minimal surfaces by considering the inverse source problem for the prescribed mean curvature equation
 $\nabla\cdot\left[\frac{\nabla u}{(1+|\nabla u|^{2})^{1/2}}\right]=H(x).$
+This work also represents the first treatment of inverse source problems for quasilinear equations.
 We prove that in two dimensions, the source function $H$ is uniquely determined by the associated Dirichlet-to-Neumann map.
 A notable feature of this problem is that although the equation is posed on an Euclidean domain, its linearization yields an anisotropic conductivity equation where the coefficient matrix corresponds to a Riemannian metric $g$ depending on the background solution.
-This work represents the first treatments of inverse source problems for quasilinear equations.
-The proof relies on the higher order linearization method.
-The main methodological contribution is the development of an approach to decouple the resulting system of nonlinear algebraic and geometric equations, which enables the complete determination of the source term from boundary measurements.
-The decoupling uses a Liouville type uniqueness result for conformal mappings.
+The main methodological contribution is the derivation of a coupled nonlinear system of algebraic and geometric partial differential equations from boundary measurements.
+Similar systems will naturally appear in other inverse problems for quasilinear equations.
+We solve the system using a Liouville type uniqueness result for conformal mappings, which recovers the source function uniquely.
 Keywords. Inverse problems, inverse source problems, prescribed mean curvature equation, quasilinear elliptic equations.
@@
 By linearizing one can then use methods for linear equations to solve the problem.
-In recent years the interest for these has grown and it started with the works [KLU18, FO20a, LLLS21a].
+In recent years the interest for these has grown and it started with the works [KLU18, FO20, LLLS21a].
 In these works the method of higher order linearization was introduced.
@@
 A map between two Riemannian manifolds is harmonic if it satisfies a certain quasilinear equation.
-Inverse problems for elliptic second order semilinear equations have been considered in [IS94, IN95, Sun10, IY13, LLLS21a, LLLS21b, LLST22, KU20, FO20b, FLL23, ST23, Nur23a, JNS23, JNS25] and inverse problems for nonlinear magnetic Schrödinger equations in [KU23, LZ23, Ma20, KUY24].
+Inverse problems for elliptic second order semilinear equations have been considered in [IS94, IN95, Sun10, IY13, LLLS21a, LLLS21b, LLST22, KU20, FO20, FLL23, ST23, Nur23a, JNS23, JNS25] and inverse problems for nonlinear magnetic Schrödinger equations in [KU23, LZ23, Ma20, KUY24].
 We also mention the works for nonlinear biharmonic equations [BKSU25, NS25], $p$-harmonic equation [BKS15, BIK18, CF25] and the works for nonlinear fractional equations [Lin22, LL19a, LL23, LL19b].
@@
 (2.15) $\lvert\nabla u_{0}\rvert=\lvert\nabla\tilde{u}_{0}\rvert|_{\phi}.$
+Remark 2.3.
+We note that prescribed Jacobian equations such as (2.14) may lack unique solutions [CDK11, Section 1.4].
+To resolve this ambiguity, we proceed to the second linearization in Section 4, which yields additional constraints on $\phi$.
+Ultimately, these constraints force $\phi$ to be a conformal mapping, whose uniqueness then follows from Liouville-type rigidity.
 3. Boundary information
@@
 That is $\phi$ is a conformal mapping in Euclidean space.
-Combining this with $\phi|_{\partial\Omega}=\mathrm{Id}$ a Liouville typer result (see e.g.
+Combining this with $\phi|_{\partial\Omega}=\mathrm{Id}$ a Liouville type result (see e.g.
 [Lio97, Proposition $3.$]) gives that $\phi$ is the identity in $\Omega$.
@@
 Acknowledgment.
-The authors would like to thank Niko Jokela for helpful discussion.
+The authors would like to thank Niko Jokela and Gyula Csató for helpful discussion.
 T.L.